Suppose that instead of plowing money back into lucrative ventures, Aqua Americas management is investing at an expected return on equity of 5%, which is below the return of 7.8% that investors could expect to get from comparable securities. a. Find the sustainable growth rate of dividends and earnings in these circum stances. Continue to assume a 40% plowback ratio. negative despite the positive growth rate of earnings and dividends. an attempted takeover? b. Find the new value of its investment opportunities. Explain why this value is c. If you were a corporate raider, would Aqua America be a good candidate for

Value of assets in place $15.81 + Present value of growth opportunities (PVGO 10.62 $26.43 - Total value of Aqua Americas stock
Answer #1

a. Sustainable growth rate formula is g = b * r , where b is retention ratio g = 0.4 * 5 = 2 %

b . New value cannot be calculated , since old value is not mentioned

c. No , Aqua america is not good to takeover as it is providing lower return .